# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 5.9 Collisionless Magnetic Reconnection
To assess whether the collisionless tearing instabilty grows fast enough to enable the onset of magnetic reconnection to disrupt the turbulent cascade, we will assess $\gamma_{RXN}/\omega_{nl}\sim\gamma_{RXN}/\omega$ assuming critical balance of linear and nonlinear timescales $\omega_{nl}\sim\omega$ (Goldreich & Sridhar, 1995; Mallet et al., 2015). Taking the linear wave frequency for MHD Alfvén waves $\omega=k_{\parallel}v_{A}=(k_{\parallel}\rho_{i})^{1/2}(k_{\perp}\rho_{i})^{1/2}v_{A}/\rho_{i}$ using (1) with $\alpha=1$ for the B06 scaling of $k_{\parallel}$, and substituting $v_{A\perp}=v_{A}(\delta B_{\perp}/B_{0})=v_{A}[(k_{0}\rho_{i})/(k_{\perp}\rho_{i})]^{1/4}$ using (4), we find that the growth rate in the $\Delta^{\prime}\delta_{in}\ll 1$ limit, which corresponds to the large unstable wavenumber $k$ limit by inspection of (48), can be expressed as
@@
The maximum growth rate $\gamma_{RXN}/\omega$ from (50) increases with $k_{\perp}\rho_{i}$, but also requires the limit $k_{\perp}\rho_{i}\ll 1$ for the validity of this linear tearing growth rate calculation. Therefore, to estimate quantitatively the maximum normalized tearing growth rate, we take $k_{\perp}\rho_{i}\rightarrow\mathcal{F}$, where $\mathcal{F}\ll 1$ is the maximum value of the normalized perpendicular wavenumber for which (50) remains valid; numerical simulations of collisionless magnetic reconnection will be invaluable to provide estimates of the maximum value of $\mathcal{F}$. The maximum growth rate $\gamma_{RXN}/\omega$ also occurs for the minimum unstable wavenumber $k\rho_{i}$ allowed by the range in (51). Thus, taking these two limits and requiring a collisionless tearing instability growth sufficient to disrupt the turbulent cascade, $\gamma_{RXN}/\omega\gtrsim 1$, yields the constraint
@@
Figure 9: (a) Scaling the normalized collisionless tearing instability growth rate $\gamma_{RXN}/\omega$ for a turblent plasma with $\beta_{i}=0.01$, $\tau=1$, $\mu=1836$, and $k_{0}\rho_{i}=10^{-4}$ as a function of the normalized unstable wavenumber $k\rho_{i}$. Colors indicate the solutions in the low $k$ limit (corresponding to $\Delta^{\prime}\delta_{in}\gtrsim 1$) (red) and in the high $k$ limit (corresponding to $\Delta^{\prime}\delta_{in}\ll 1$) (blue). Line thickness indicates the either a sinusoidal current profile with $n=2$ (thick) or a Harris-like hyperbolic tangent current profile with $n=1$ (thin). Line styles indicate different values of maximum perpendicular wavenumber $\mathcal{F}=1.0$ (solid), $0.3$ (dashed), and $0.1$ (dotted). (b) The corresponding normalized tearing mode instability parameter $\Delta^{\prime}\delta_{in}$ vs. $k\rho_{i}$ for each case.
@@
Once again the maximum growth rate $\gamma_{RXN}/\omega$ from (54) increases with $k_{\perp}\rho_{i}$ (althoughly weakly with the $1/4$ power), so once again we take the limit $k_{\perp}\rho_{i}\rightarrow\mathcal{F}$, where $\mathcal{F}\ll 1$ is the maximum value of the normalized perpendicular wavenumber for which (54) remains valid. The maximum growth rate will also occur for the maximum unstable wavenumber $k\rho_{i}$ given by (55). Thus, taking these two limits and requiring a collisionless tearing instability growth rate sufficient to distrupt the turbulent cascade, $\gamma_{RXN}/\omega\gtrsim 1$, yields the constraint
@@
It is necessary to highlight here a few important caveats in using the linear collisionless tearing mode growth rates and MHD turbulence scaling theory to predict when magnetic reconnection may disrupt the dynamics of the turbulent cascade. First, the linear tearing mode growth rates used here are rigorously derived in asymptotic limits, but we employ these same rates up to the boundary of those limits. For example, the linear growth rates formally require $k_{\perp}\rho_{i}\ll 1$; also, the $\Delta^{\prime}\delta_{in}\ll 1$ limit is used to estimate the tearing mode growth rates as $\Delta^{\prime}\delta_{in}\rightarrow 1$. But such rates derived in asymptotic limits may not remain quantitatively accurate as we extend beyond their regime of formal validity. We take the approach that the growth rates from the limits are at least a reasonable lowest-order estimate of the rate beyond their formally limited range. Furhermore, the growth rates for the collisionless tearing mode are determined under smooth conditions, whereas to determine whether reconnection arises in the current sheets that naturally arise in plasma turbulence one needs to determine the collisionless tearing growth rates in the presence of turbulence; once again, the estimated tearing mode growth rates used here we take simply as the lowest-order estimation. Computing the tearing mode growth rates in a turbulent plasma, and determining the competition of the tearing mode with the nonlinear cascade of energy due to turbulence, are both necessary for more accurate calculations, which likely will require detailed kinetic numerical simulations. Finally, alternative proposed mechanisms for the development of current sheets in plasma turbulence—such as the nonlinear interaction of counterpropagating Alfvén wave collisions (Howes, 2016), as opposed to dynamic alignment (Boldyrev, 2006)—may yield a different scaling for the width and thickness of current sheets; the resulting predictions for the onset of magnetic reconnection in a turbulent plasma would thereby also likely change.
@@
In summary, we have followed previous investigations (Loureiro & Boldyrev, 2017a; Mallet et al., 2017b) to estimate how the collisionless tearing mode may arise in the current sheets that develop self-consistently in plasma turbulence, potentially disrupting the nonlinear cascade of turbulent energy to small scales. We restrict ourselves to the moderately low electron plasma beta limit $\mu^{-1}\ll\beta_{e}\ll 1$, which can be expressed in terms of the turbulence parameters in Table 3 as the limit $1\ll\beta_{i}\mu/\tau\ll\mu$. In this limit, magnetic flux conservation is broken due to electron inertia at electron inertial length scales $d_{e}$ (Zocco & Schekochihin, 2011; Loureiro & Boldyrev, 2017a; Mallet et al., 2017b). We consider only the case when collisionless magnetic reconnection arises in current sheets with thicknesses $a\sim 1/k_{\perp}$ in the MHD regime $k_{\perp}\rho_{i}\lesssim 1$. We find that the growth rate of reconnection relative to the turbulent fluctuation frequencies $\gamma_{RXN}/\omega$ is a function of the isotropic driving wavenumber $k_{0}\rho_{i}$, ion plasma beta $\beta_{i}$, and ion-to-electron temperature ratio $\tau$, as well as the ion-to-electron mass ratio $\mu$.
@@
In general, the normalized collisionless tearing instabilty growth rates $\gamma_{RXN}/\omega$—given by (50) in the $\Delta^{\prime}\delta_{in}\ll 1$ limit and by (54) in the $\Delta^{\prime}\delta_{in}\gtrsim 1$ limit—increase as $k_{0}\rho_{i}$ and $\beta_{i}$ decrease. For specified values of the plasma parameters $\beta_{i}$ and $\tau$, we obtain different conditions on the maxmimum value of $k_{0}\rho_{i}$ needed for reconnection to arise, given by (52) and (53) in the $\Delta^{\prime}\delta_{in}\ll 1$ limit and by (57) and (58) in the $\Delta^{\prime}\delta_{in}\gtrsim 1$ limit. These scalings mean that reconnection is more likely to arise in turbulence with a very large driving scale relative to the ion Larmor radius scale $k_{0}\rho_{i}\ll 1$ and for low ion plasma beta condtions $\beta_{i}\ll 1$.
@@
It is worthwhile emphasizing that these estimates on the importance of magnetic reconnection in plasma turbulence are calculated only if for current sheets with a thickness $a\sim 1/k_{\perp}$ in the MHD regime $k_{\perp}\rho_{i}\ll 1$ so that the MHD scaling predictions in §3 apply. It is also possible that electron-only magnetic reconnection (Phan et al., 2018) could occur with the range of scales in the kinetic dissipation range $k_{\perp}\rho_{i}\ll 1$, but to estimate whether this will occur would require the use of modified turbulence scalings appropriate for the dissipation range (Howes et al., 2008a, 2011).
@@
Finally, how the incorporation of magnetic reconnection into the turbulent cascade impacts the damping of the turbulence and the resulting energization of the plasma remains an open question. Collisionless magnetic reconnection effectively converts magnetic energy into other forms, specifically the kinetic energy of the bulk plasma outflows and possibly nonthermal energy in the particle velocity distribution functions (Drake et al., 2005; Egedal et al., 2008, 2012; Drake & Swisdak, 2012; Loureiro et al., 2013; Jiansen et al., 2018; Pucci et al., 2018; Muñoz & Büchner, 2018; Dahlin, 2020; McCubbin et al., 2022). But, the linear terms in the equations of evolution for the turbulent plasma mediate the ongoing oscillation of energy from magnetic energy to the kinetic energy of the turbulent bulk plasma motions and back (Howes, 2015). For example, the physics of undamped Alfvén waves involve the transformation of the bulk kinetic energy of the perpendicular wave motions to magnetic energy as the frozen-in magnetic field is stretched; magnetic tension serves as the restoring force for the wave, decelerating the perpendicular wave motions and ultimately re-accelerating them back towards the state where the magnetic field is not bent, thereby converting magnetic energy back to kinetic energy. Thus, collisionless magnetic reconnection does not necessarily lead directly to damping of the turbulent motions, but rather is simply a channel, in additon to magnetic tension, for converting magnetic to kinetic energy. The triggering of magnetic reconnection, however, will impact the scaling of the three-dimensional wave vector of the turbulent motions that result from the reconnection flow (Boldyrev & Loureiro, 2017; Loureiro & Boldyrev, 2017b, a; Mallet et al., 2017b, a; Walker et al., 2018), and these different scalings may channel energy into different kinetic physical mechanisms that transfer turbulent energy into particle energy.
@@ 6.1 Collisionality of the Dissipation Range of Plasma Turbulence
Figure 10: Phase Diagram of the extent of the turbulent inertial range from driving scales at $k_{\parallel 0}\lambda_{mfp,e}$ to the transition to the dissipation range at $k_{\parallel T}\lambda_{mfp,e}$ vs.  the ion plasma beta $\beta_{i}$. The plot indicates that, at the small scale end of the inertial range where damping mechanisms can remove energy from the turbulence, for many plasma systems of interest, the governing dynamics are weakly collisional with t $k_{\parallel T}\lambda_{mfp,e}\gg 1$, so kinetic plasma theory is necessary to capture the physics of turbulent dissipation.
@@
We first create a phase diagram indicating that, for many plasmas of interest, even if the turbulence is driven at fluid scales with $k_{\parallel,0}\lambda_{mfp,e}\ll 1$, the end of the inertial range generally occurs at scales that are weakly collisional. Therefore, the kinetic damping mechanisms analyzed here are responsible for removing energy from the turbulent cascade, rather than viscosity and resistivity, which are based on microscopic Coulomb collisions between particles (Braginskii, 1965; Chapman & Cowling, 1970). Since the mean free path must be compared to the parallel length scales of the turbulent fluctuations characterized by $k_{\parallel}$, we must first determine the parallel length scales associated with the small-scale end of the inertial range at the transition to the dissipation range, corresponding to a perpendicular scale $k_{\perp}\rho_{i}\sim 1$. Specifying the GS95 scaling with $\alpha=0$ for this example, the parallel wavenumber at the transiton to the dissipation range $k_{\parallel T}$, determined by (1), is given by $k_{\parallel T}\rho_{i}=(k_{0}\rho_{i})^{1/3}(k_{\perp}\rho_{i})^{2/3}=(k_{0}\rho_{i})^{1/3}$, here taking $k_{\perp}\rho_{i}\sim 1$ as the end of the inertial range. For simplicity, we assume here that the turbulence is driven strongly and isotropically with $\chi_{0}=1$ and $k_{\parallel 0}/k_{\perp 0}=1$. Normalizing this parallel scale in terms of the electron collisional mean free path $\lambda_{mfp,e}$, we obtain $k_{\parallel T}\lambda_{mfp,e}=(k_{0}\rho_{i})^{1/3}(\lambda_{mfp,e}/\rho_{i})$.
@@
The range of plasma and turbulence parameters for a number of plasma systems of interest are denoted by the colored boxes in Fig. 10. Plotted on the horizontal axis of Fig. 10 is the range of parallel scales bounding the turbulent inertial range normalized to the electron collisional mean free path $\lambda_{mfp,e}$: (i) the normalized parallel driving scale $k_{\parallel 0}\lambda_{mfp,e}$ of the turbulence is represented by the left end of each box; (ii) the turbulent energy cascades to smaller scales to the right within each box (wavy red arrow below plot); and (iii) the parallel scale corresponding to the small-scale end of the inertial range at $k_{\parallel T}\lambda_{mfp,e}$ is represented by the right end of each box. The vertical extent of the colored boxes indicates the range of ion plasma beta $\beta_{i}$ for each of the chosen systems. The fluid regime (yellow region) indicates conditions under which the dynamics of the turbulent fluctuations occur on parallel scales that are collisional with $k_{\parallel 0}\lambda_{mfp,e}\ll 1$; the kinetic regime (light blue region) indicates conditions under which the dynamics of the turbulent fluctuations occur on parallel scales that are weakly collisional with $k_{\parallel 0}\lambda_{mfp,e}\gg 1$.
@@
In Fig. 10 are shown these parameter ranges for the near Earth solar wind (red), the Earth’s magnetosheath and the solar corona (green), the inner helisophere at heliocentric radii $R\ll 1$ AU and the outer helisophere at $R\gg 1$ AU (blue), the Galactic Center and intracluster medium (black), and terrestrial laboratory experiments including the Large Plasma Device (LAPD)at UCLA (Gekelman et al., 1991, 2016) and the Joint European Torus (JET) Fusion experiment (JET Team, 1992). The main message of Fig. 10 is that, for most turbulent plasma systems of interest, the parallel length scales associated with the end of the inertial range correspond to weakly collisional plasma conditions with $k_{\parallel T}\lambda_{mfp,e}\gg 1$. Therefore, to explore the physical damping mechanisms responsible for removing energy from the turbulent fluctuations and consequently energizing the plasma particles, plasma kinetic theory is essential.

### v2

@@ 5.9 Collisionless Magnetic Reconnection
To assess whether the collisionless tearing instability grows fast enough to enable the onset of magnetic reconnection to disrupt the turbulent cascade, we will assess $\gamma_{RXN}/\omega_{nl}\sim\gamma_{RXN}/\omega$, assuming critical balance of linear and nonlinear timescales $\omega_{nl}\sim\omega$ (Goldreich & Sridhar, 1995; Mallet et al., 2015). Taking the linear wave frequency for MHD Alfvén waves $\omega=k_{\parallel}v_{A}=(k_{\parallel}\rho_{i})^{1/2}(k_{\perp}\rho_{i})^{1/2}v_{A}/\rho_{i}$ using (1) with $\alpha=1$ for the B06 scaling of $k_{\parallel}$, and substituting $v_{A\perp}=v_{A}(\delta B_{\perp}/B_{0})=v_{A}[(k_{0}\rho_{i})/(k_{\perp}\rho_{i})]^{1/4}$ using (4), we find that the growth rate in the $\Delta^{\prime}\delta_{in}\ll 1$ limit, which corresponds to the large unstable wavenumber $k$ limit by inspection of (52), can be expressed as
@@
The maximum growth rate $\gamma_{RXN}/\omega$ from (54) increases with $k_{\perp}\rho_{i}$, but also requires the limit $k_{\perp}\rho_{i}\ll 1$ for the validity of this linear tearing growth rate calculation. Therefore, to estimate quantitatively the maximum normalized tearing growth rate, we take $k_{\perp}\rho_{i}\rightarrow\mathcal{F}$, where $\mathcal{F}\ll 1$ is the maximum value of the normalized perpendicular wavenumber for which (54) remains valid; numerical simulations of collisionless magnetic reconnection will be invaluable to provide estimates of the maximum value of $\mathcal{F}$. The maximum growth rate $\gamma_{RXN}/\omega$ also occurs for the minimum unstable wavenumber $k\rho_{i}$ allowed by the range in (55). Thus, taking these two limits and requiring a collisionless tearing instability growth rate sufficient to disrupt the turbulent cascade, $\gamma_{RXN}/\omega\gtrsim 1$, yields the constraint
@@
Figure 10: (a) Scaling the normalized collisionless tearing instability growth rate $\gamma_{RXN}/\omega$ for a turbulent plasma with $\beta_{i}=0.01$, $\tau=1$, $\mu=1836$, and $k_{0}\rho_{i}=10^{-4}$ as a function of the normalized unstable wavenumber $k\rho_{i}$. Colors indicate the solutions in the low $k$ limit (corresponding to $\Delta^{\prime}\delta_{in}\gtrsim 1$) (red) and in the high $k$ limit (corresponding to $\Delta^{\prime}\delta_{in}\ll 1$) (blue). Line thickness indicates the either a sinusoidal current profile with $n=2$ (thick) or a Harris-like hyperbolic tangent current profile with $n=1$ (thin). Line styles indicate different values of maximum perpendicular wavenumber $\mathcal{F}=1.0$ (solid), $0.3$ (dashed), and $0.1$ (dotted). (b) The corresponding normalized tearing mode instability parameter $\Delta^{\prime}\delta_{in}$ vs. $k\rho_{i}$ for each case.
@@
Once again the maximum growth rate $\gamma_{RXN}/\omega$ from (58) increases with $k_{\perp}\rho_{i}$ (although weakly with the $1/4$ power), so once again we take the limit $k_{\perp}\rho_{i}\rightarrow\mathcal{F}$, where $\mathcal{F}\ll 1$ is the maximum value of the normalized perpendicular wavenumber for which (58) remains valid. The maximum growth rate will also occur for the maximum unstable wavenumber $k\rho_{i}$ given by (59). Thus, taking these two limits and requiring a collisionless tearing instability growth rate sufficient to disrupt the turbulent cascade, $\gamma_{RXN}/\omega\gtrsim 1$, yields the constraint
@@
It is necessary to highlight here a few important caveats in using the linear collisionless tearing mode growth rates and MHD turbulence scaling theory to predict when magnetic reconnection may disrupt the dynamics of the turbulent cascade. First, the linear tearing mode growth rates used here are rigorously derived in asymptotic limits, but we employ these same rates up to the extreme boundary of those limits. For example, the linear growth rates formally require $k_{\perp}\rho_{i}\ll 1$, but we take $k_{\perp}\rho_{i}\rightarrow 1$ to estimate the maximum growth rate; also, the $\Delta^{\prime}\delta_{in}\ll 1$ limit is used to estimate the tearing mode growth rates as $\Delta^{\prime}\delta_{in}\rightarrow 1$. Such rates derived in asymptotic limits, however, may not remain quantitatively accurate as we extend beyond their formal regime of validity. We adopt the viewpoint that the growth rates calculated in these limits are at least reasonable lowest-order estimates of the rates even beyond their formally limited range. Furthermore, the growth rates for the collisionless tearing mode are determined under smooth conditions, whereas to determine whether reconnection arises in the current sheets that naturally arise in plasma turbulence, one needs to determine the collisionless tearing growth rates in the presence of turbulence; once again, the estimated tearing mode growth rates used here are taken simply as the lowest-order estimation. Computing the tearing mode growth rates in a turbulent plasma and determining the competition of the tearing mode with the nonlinear cascade of energy due to turbulence are both necessary for more accurate calculations, which likely will require detailed kinetic numerical simulations. Finally, alternative proposed mechanisms for the development of current sheets in plasma turbulence—such as the nonlinear interaction of counterpropagating Alfvén wave collisions (Howes, 2016), as opposed to dynamic alignment (Boldyrev, 2006)—may yield a different scaling for the width and thickness of current sheets; the resulting predictions for the onset of magnetic reconnection in a turbulent plasma would thereby also likely change.
@@
In summary, we have followed previous investigations (Loureiro & Boldyrev, 2017a; Mallet et al., 2017b) to estimate how the collisionless tearing mode may arise in the current sheets that develop self-consistently in plasma turbulence, potentially disrupting the nonlinear cascade of turbulent energy to small scales. We restrict ourselves to the moderately low electron plasma beta limit $\mu^{-1}\ll\beta_{e}\ll 1$, which can be expressed in terms of the plasma and turbulence parameters in Table 4 as the limit $1\ll\beta_{i}\mu/\tau\ll\mu$. In this limit, magnetic flux conservation is broken due to electron inertia at electron inertial length scales $d_{e}$ (Zocco & Schekochihin, 2011; Loureiro & Boldyrev, 2017a; Mallet et al., 2017b). We consider only the case when collisionless magnetic reconnection arises in current sheets with thicknesses $a\sim 1/k_{\perp}$ in the MHD regime $k_{\perp}\rho_{i}\lesssim 1$. We find that the growth rate of reconnection relative to the turbulent fluctuation frequencies $\gamma_{RXN}/\omega$ is a function of the isotropic driving wavenumber $k_{0}\rho_{i}$, ion plasma beta $\beta_{i}$, and ion-to-electron temperature ratio $\tau$, as well as the ion-to-electron mass ratio $\mu$.
@@
In general, the normalized collisionless tearing instability growth rates $\gamma_{RXN}/\omega$—given by (54) in the $\Delta^{\prime}\delta_{in}\ll 1$ limit and by (58) in the $\Delta^{\prime}\delta_{in}\gtrsim 1$ limit—increase as both $k_{0}\rho_{i}$ and $\beta_{i}$ decrease. For specified values of the plasma parameters $\beta_{i}$ and $\tau$, we obtain different conditions on how small $k_{0}\rho_{i}$ must be for reconnection to arise, given by (56) and (57) in the $\Delta^{\prime}\delta_{in}\ll 1$ limit and by (61) and (62) in the $\Delta^{\prime}\delta_{in}\gtrsim 1$ limit. These scalings mean that reconnection is more likely to arise in turbulence with a very large driving scale relative to the ion Larmor radius scale—meaning a smaller value of $k_{0}\rho_{i}\ll 1$—and for low ion plasma beta conditions $\beta_{i}\ll 1$.
@@
It is worthwhile emphasizing that these estimates on the importance of magnetic reconnection in plasma turbulence are calculated only for current sheets with a thickness $a\sim 1/k_{\perp}$ in the MHD regime $k_{\perp}\rho_{i}\ll 1$ so that the MHD scaling predictions in §3 apply. It is also possible that electron-only magnetic reconnection (Phan et al., 2018) could occur within the range of scales in the kinetic dissipation range $k_{\perp}\rho_{i}\gtrsim 1$, but to estimate whether this will occur would require the use of modified turbulence scalings appropriate for the dissipation range (Howes et al., 2008a, 2011).
@@
Finally, how the incorporation of magnetic reconnection into the turbulent cascade impacts the damping of the turbulence and the resulting energization of the plasma remains an open question. Collisionless magnetic reconnection effectively converts magnetic energy into other forms, specifically the kinetic energy of the bulk plasma outflows and possibly internal energy in the particle velocity distribution functions (Drake et al., 2005; Egedal et al., 2008, 2012; Drake & Swisdak, 2012; Loureiro et al., 2013; Jiansen et al., 2018; Pucci et al., 2018; Muñoz & Büchner, 2018; Dahlin, 2020; McCubbin et al., 2022). But such conversion between magnetic energy and the bulk kinetic energy of turbulent plasma motions already occurs ubiquitously in plasma turbulence: the linear terms in the equations of evolution for the turbulent plasma mediate the ongoing oscillation of energy from magnetic energy to the kinetic energy of the turbulent bulk plasma motions and back (Howes, 2015). For example, the physics of undamped Alfvén waves involve the transformation of the bulk kinetic energy of the perpendicular wave motions to magnetic energy as the frozen-in magnetic field is stretched; magnetic tension serves as the restoring force for the wave, decelerating the perpendicular wave motions and ultimately re-accelerating them back towards the state where the magnetic field is not bent, thereby converting magnetic energy back to kinetic energy. Thus, collisionless magnetic reconnection does not necessarily lead directly to damping of the turbulent motions (unless a significant fraction of the released magnetic energy is converted into internal energy of the particle velocity distributions), but rather is simply a channel, in addition to magnetic tension, for converting magnetic to kinetic energy. The triggering of magnetic reconnection, however, will impact the scaling of the characteristic three-dimensional wave vector of the turbulent motions that result from the reconnection flow (Boldyrev & Loureiro, 2017; Loureiro & Boldyrev, 2017b, a; Mallet et al., 2017b, a; Walker et al., 2018), and these different scalings may channel energy through different kinetic physical mechanisms that ultimately transfer turbulent energy to internal particle energy. Thus, the partitioning of turbulent energy between ions and electrons, $Q_{i}/Q_{e}$, and between degrees of freedom for both species, $Q_{\perp,i}/Q_{\parallel,i}$ and $Q_{\perp,e}/Q_{\parallel,e}$, remains to be ascertained in situations where collisionless magnetic reconnection is triggered in plasma turbulence.
@@ 6.1 Collisionality of the Dissipation Range of Plasma Turbulence
Figure 11: Phase Diagram of the extent of the turbulent inertial range from driving scales at $k_{\parallel 0}\lambda_{mfp,e}$ to the transition to the dissipation range at $k_{\parallel T}\lambda_{mfp,e}$ vs. the ion plasma beta $\beta_{i}$. The plot indicates that, at the small scale end of the inertial range where damping mechanisms can remove energy from the turbulence, for many plasma systems of interest the governing dynamics are weakly collisional with $k_{\parallel T}\lambda_{mfp,e}\gg 1$, so kinetic plasma theory is necessary to capture the physics of turbulent dissipation.
@@
We first create a phase diagram indicating that, for many plasmas of interest, even if the turbulence is driven at fluid scales with $k_{\parallel,0}\lambda_{mfp,e}\ll 1$, the end of the inertial range generally occurs at scales that are weakly collisional. Therefore, the kinetic damping mechanisms analyzed here are responsible for removing energy from the turbulent cascade, rather than viscosity and resistivity, which are based on microscopic Coulomb collisions between particles in the limit of strong collisionality (Braginskii, 1965; Chapman & Cowling, 1970). Since the mean free path must be compared to the parallel length scales of the turbulent fluctuations characterized by $k_{\parallel}$, we must first determine the parallel length scales associated with the small-scale end of the inertial range at the transition to the dissipation range, corresponding to a perpendicular scale $k_{\perp}\rho_{i}\sim 1$. Specifying the GS95 scaling with $\alpha=0$ for this example, the parallel wavenumber at the transition to the dissipation range $k_{\parallel T}$, determined by (1), is given by $k_{\parallel T}\rho_{i}=(k_{0}\rho_{i})^{1/3}(k_{\perp}\rho_{i})^{2/3}=(k_{0}\rho_{i})^{1/3}$, here taking $k_{\perp}\rho_{i}\sim 1$ as the end of the inertial range. For simplicity, we assume here that the turbulence is driven strongly and isotropically with $\chi_{0}=1$ and $k_{\parallel 0}/k_{\perp 0}=1$. Normalizing this parallel scale in terms of the electron collisional mean free path $\lambda_{mfp,e}$, we obtain $k_{\parallel T}\lambda_{mfp,e}=(k_{0}\rho_{i})^{1/3}(\lambda_{mfp,e}/\rho_{i})$.
@@
The range of plasma and turbulence parameters for a number of plasma systems of interest are denoted by the colored boxes in Fig. 11. Plotted on the horizontal axis of Fig. 11 is the range of parallel scales bounding the turbulent inertial range normalized to the electron collisional mean free path $\lambda_{mfp,e}$: (i) the normalized parallel driving scale $k_{\parallel 0}\lambda_{mfp,e}$ of the turbulence is represented by the left end of each box; (ii) the turbulent energy cascades to smaller scales to the right within each box (wavy red arrow below plot); and (iii) the parallel scale corresponding to the small-scale end of the inertial range at $k_{\parallel T}\lambda_{mfp,e}$ is represented by the right end of each box. The vertical extent of the colored boxes indicates the range of ion plasma beta $\beta_{i}$ for each of the chosen systems. The fluid regime (yellow region) indicates conditions under which the dynamics of the turbulent fluctuations occur on parallel scales that are strongly collisional with $k_{\parallel 0}\lambda_{mfp,e}\ll 1$; the kinetic regime (light blue region) indicates conditions under which the dynamics of the turbulent fluctuations occur on parallel scales that are weakly collisional with $k_{\parallel 0}\lambda_{mfp,e}\gg 1$.
@@
In Fig. 11 are shown these parameter ranges for a variety of plasma systems: the near Earth solar wind (red); the Earth’s magnetosheath and the solar corona (green); the inner heliosphere at heliocentric radii $R\ll 1$ AU and the outer heliosphere at $R\gg 1$ AU (blue); the interstellar medium, the Galactic Center, and the intracluster medium (black),; and terrestrial laboratory experiments including the Large Plasma Device (LAPD) at UCLA (Gekelman et al., 1991, 2016) and the Joint European Torus (JET) Fusion experiment (JET Team, 1992). The main message of Fig. 11 is that, for most turbulent plasma systems of interest, the parallel length scales associated with the end of the inertial range correspond to weakly collisional plasma conditions with $k_{\parallel T}\lambda_{mfp,e}\gg 1$. Therefore, to explore the physical damping mechanisms responsible for removing energy from the turbulent fluctuations and consequently energizing the plasma particles, plasma kinetic theory is essential.